OpenStudy (anonymous):

What is the solution to the linear system of equations? A. (2, –3) B. (5, 0) C. The system has no solution. D. (2, 3)

OpenStudy (sidsiddhartha):

\[3x+y=3\] \[x-y=5\] u can start with elimination anything x or y so if u eliminate x then from the second eq. u can write \[x=y+5\] now put this in the first eq. in place of x \[3(y+5)+y=3\] now solve for y

OpenStudy (texaschic101):

nicely explained :)

OpenStudy (sidsiddhartha):

thank u @texaschic101

OpenStudy (texaschic101):

just telling the truth....I give credit where credit is due :)

OpenStudy (sidsiddhartha):

:)

OpenStudy (sidsiddhartha):

hay @Grim4 now can u solve it?

OpenStudy (anonymous):

y=-3

OpenStudy (sidsiddhartha):

yeah y=-3 now can u find out x

OpenStudy (texaschic101):

3(y + 5) + y = 3 3y + 15 + y = 3 3y + y = 3 - 15 4y = - 12 y = -3 yep...now sub -3 back into either equation to find x

OpenStudy (anonymous):

x= -8

OpenStudy (anonymous):

Now that I think about it i think that might be wrong o_O

OpenStudy (sidsiddhartha):

remember the second equation \[x-y=5\]so \[x=y+5\] now put y=-3

OpenStudy (anonymous):

It's 2

OpenStudy (sidsiddhartha):

yup correct :)

OpenStudy (anonymous):

Thanks for helping

OpenStudy (sidsiddhartha):

np :)